### v1

@@ 5. Numerical simulations
In this Section, we will now present an Algorithm for the solution of (1.2) in case of a bounded linear operator $T:\mathbf{L}^{2}\left(\Omega\right)\to\mathbf{L}^{2}\left(\mathbb{M}\right)$ with $\mathcal{R}(u)=\left\|u\right\|_{\mathbf{L}^{2}\left(\Omega\right)}^{2}$, i.e.
@@
Note, that we have set $F_{h}=F=T$ here. In this linear setup with quadratic penalty, existence and uniqueness of a minimizer can readily be derived as in [SGG+08].
@@ 5.1. Numerical algorithm
In [HW14], the problem (5.1) has been solved by means of the dual problem, corresponding to a quadratic minimization problem under box constraints, which was then solved using the quadprog algorithm. This method has several advantages, including computable bounds for the accuracy by means of the primal-dual gap, but relying on quadprog makes it slow and especially not applicable for large-scale problems, where the matrix corresponding to $F_{h}$ cannot be stored or set up explicitly.
@@
Note, that (ADMM1) can be solved explicitly by means of the necessary and sufficient first-order optimality condition, yielding
@@
where the application is meant point-wise.
@@
Note, that this can also be interpreted as linearizing the quadratic term in (ADMM1) and adding a penalty, which is why this method is called the alternating direction linearized proximal method of multipliers (AD-LPMM). The resulting algorithm is as follows:
@@ 5.2. Test problem
Note, that $\left(Tu\right)^{\prime\prime}=-u$ for any function $u\in\mathbf{L}^{2}\left(\left[0,1\right]\right)$, so that estimating $u$ from $Tu$ corresponds to estimating the second derivative of $Tu$. Obviously, $T$ is $2$-smoothing, such that (2.3) and (2.8) are both satisfied with $a=2$.
@@ 5.3. Results
To simulate the rate of convergence, we use $n=257$ in (5.2) and $M=100$ Monte Carlo runs. In each run, we simulate data $g^{\mathrm{obs}}$ as in (4.1) and perform (5.1) both with ADMM and AD-LPMM as well as (5.3). Therein, we perform $K=10.000$ iterations both in ADMM and AD-LPMM. The regularization parameter $\alpha$ is varied over $40$ different values, which are chosen logarithmically equispaced between $10^{-8}$ and $10^{-2}$, separately for each set of data. Afterwards, we determine that value of $\alpha$, which leads the minimal error $\frac{1}{M}\sum_{m=1}^{M}\left\|\hat{u}_{\alpha}^{i}-u^{\dagger}\right\|_{\mathbf{L}^{2}}^{2}$, where $i$ denotes the reconstruction in the $i$th run, i=1,,Mi=1,...,M. This way, we simulate the optimal $\alpha$ as in Theorem 4.6. The above procedure is repeated for different values of $\sigma$, and the corresponding optimal error is plotted against $\sigma$ in Figure 1 below.
@@ 6. Conclusion
In this paper, we have investigated $\mathbf{L}^{1}$ data fitting for Inverse Problems with discretized noises. We have derived error bounds in a general setting, including noise in the forward operator, under verifiable conditions. Applied to the situation of $a$-smoothing operators and Gaussian white noise, this yields order-optimal rates of convergence. Therefore, the usage of $\mathbf{L}^{1}$ data fitting comes at no price, at least asymptotically. We have also discussed possible algorithms for the solution of the minimization problem, which allow to compute the corresponding minimizers even in large-scale real world applications. In view of the increased robustness of $\mathbf{L}^{1}$ data fitting compared to standard $\mathbf{L}^{2}$ based approaches, it seems therefore reasonable to use $\mathbf{L}^{1}$ data fitting more often.
In interesting question for future research, which was not treated in this paper, is how to choose the parameter $\alpha$ in (1.2). As discussed in Section 4, the derived order-optimal convergence rates are only valid for an optimal choice of $\alpha$ depending on the thresholding parameter $\lambda$, which itself is chosen depending on unknown quanitites. Therefore, an a posteriori parameter choice rule would be of immediate interest.

### v2

@@ 5. Numerical simulations
In this section, we will now present an Algorithm for the solution of (1.2) in case of a bounded linear operator $T:\mathbf{L}^{2}\left(\Omega\right)\to\mathbf{L}^{1}\left(\mathbb{M}\right)$ with $\mathcal{R}(u)=\left\|u\right\|_{\mathbf{L}^{2}\left(\Omega\right)}^{2}$, i.e.
@@
Note that we have set $F_{h}=F=T$ here. In this linear setup with quadratic penalty, existence and uniqueness of a minimizer can readily be derived as in [SGG+08].
@@ 5.1. Numerical algorithm
In [HW14], the problem (5.1) has been solved by means of the dual problem, corresponding to a quadratic minimization problem under box constraints, which was then solved using the quadprog algorithm. This method has several advantages, including computable bounds for the accuracy by means of the primal-dual gap, but relying on quadprog makes it slow and is especially not applicable for large-scale problems, where the matrix corresponding to $F_{h}$ cannot be stored or set up explicitly.
@@
Note that (ADMM1) can be solved explicitly by means of the necessary and sufficient first-order optimality condition, yielding
@@
where the application is performed point-wise.
@@
Note that this can also be interpreted as linearizing the quadratic term in (ADMM1) and adding a penalty, which is why this method is called the alternating direction linearized proximal method of multipliers (AD-LPMM). The resulting algorithm is as follows:
@@ 5.2. Test problem
Note that $\left(Tu\right)^{\prime\prime}=-u$ for any function $u\in\mathbf{L}^{2}\left(\left[0,1\right]\right)$, so that estimating $u$ from $Tu$ corresponds to estimating the second derivative of $Tu$. Obviously, $T$ is $2$-smoothing, such that (2.3) and (2.8) are both satisfied with $a=2$. Since $a>d/2=1/2$, we can also consider $T$ as an operator mapping into $\mathbf{L}^{1}\left([0,1]\right)$.
@@ 5.3. Results
To simulate the rate of convergence, we use $n=257$ in (5.2) and $M=100$ Monte Carlo runs. In each run, we simulate data $g^{\mathrm{obs}}$ as in (4.1) and perform (5.1) both with ADMM and AD-LPMM as well as (5.3). Therein, we perform $K=10,000$ iterations both in ADMM and AD-LPMM. The regularization parameter $\alpha$ is varied over $40$ different values, which are chosen logarithmically equispaced between $10^{-8}$ and $10^{-2}$, separately for each set of data. Afterwards, we determine that value of $\alpha$, which leads to the minimal error $\frac{1}{M}\sum_{m=1}^{M}\left\|\hat{u}_{\alpha}^{i}-u^{\dagger}\right\|_{\mathbf{L}^{2}}^{2}$, where $i$ denotes the reconstruction in the $i$th run, i=1,,Mi=1,...,M. This way, we simulate the optimal $\alpha$ as in Theorem 4.6. The above procedure is repeated for different values of $\sigma$, and the corresponding optimal error is plotted against $\sigma$ in Figure 1 below.
@@ 6. Conclusion
In this paper, we have investigated $\mathbf{L}^{1}$ data fitting for Inverse Problems with discretized noises. We have derived error bounds in a general setting, including noise in the forward operator, under verifiable conditions. Applied to the situation of $a$-smoothing operators and Gaussian white noise, this yields order-optimal rates of convergence. Therefore, the usage of $\mathbf{L}^{1}$ data fitting comes at no price, at least asymptotically. We have also discussed possible algorithms for the solution of the minimization problem, which allow computing the corresponding minimizers even in large-scale real world applications. In view of the increased robustness of $\mathbf{L}^{1}$ data fitting compared to standard $\mathbf{L}^{2}$ based approaches, it seems therefore reasonable to use $\mathbf{L}^{1}$ data fitting more often.
An interesting question for future research, which was not treated in this paper, is how to choose the parameter $\alpha$ in (1.2). As discussed in Section 4, the derived order-optimal convergence rates are only valid for an optimal choice of $\alpha$ depending on the thresholding parameter $\lambda$, which itself is chosen depending on unknown quantities. Therefore, an a posteriori parameter choice rule would be of immediate interest.